# Exploring Multilingual Human Value Concepts in Large Language Models: Is Value Alignment Consistent, Transferable and Controllable across Languages?

### v1

@@ Abstract
Prior research in representation engineering has revealed that LLMs encode concepts within their representation spaces, predominantly centered around English. In this study, we extend this philosophy to a multilingual scenario, delving into multilingual human value concepts in LLMs. Through our comprehensive exploration covering 7 types of human values, 16 languages and 3 LLM series with distinct multilinguality, we empirically substantiate the existence of multilingual human values in LLMs. Further cross-lingual analysis on these concepts discloses 3 traits arising from language resource disparities: cross-lingual inconsistency, distorted linguistic relationships, and unidirectional cross-lingual transfer between high- and low-resource languages, all in terms of human value concepts. Additionally, we validate the feasibility of cross-lingual control over value alignment capabilities of LLMs, leveraging the dominant language as a source language. Drawing from our findings on multilingual value alignment, we prudently provide suggestions on the composition of multilingual data for LLMs pre-training: including a limited number of dominant languages for cross-lingual alignment transfer while avoiding their excessive prevalence, and keeping a balanced distribution of non-dominant languages. We aspire that our findings would contribute to enhancing the safety and utility of multilingual AI.
@@
Exploring Multilingual Human Value Concepts in Large Language Models: Is Value Alignment Consistent, Transferable and Controllable across Languages?
[…]
@@ 1 Introduction
Recent years have witnessed the emergence of large language models, such as ChatGPT (OpenAI, 2023a), GPT-4 (OpenAI, 2023b), and LLaMA2 (Touvron et al., 2023). These LLMs have shown powerful capabilities in natural language understanding and generation (Guo et al., 2023; Bang et al., 2023; Jiao et al., 2023). However, alongside with their prowess, LLMs present potential threats to humanity. Research has demonstrated that LLMs can generate responses containing toxic, untruthful, biased, and even illegal content (Cui et al., 2024; Wang et al., 2023; Huang et al., 2023). Thus, aligning LLMs with human values (i.e., value alignment) is necessary for unleashing their potential safely.
@@
Human values, encompassing concepts like fairness, morality, utilitarianism, and so on, although challenging to be precisely defined in language, are undoubtedly embedded in textual form (Hendrycks et al., 2021). Recent studies in representation engineering (Zou et al., 2023a) have unveiled that LLMs encode representations of these concepts. They utilize positive and negative text pairs, aligned with the directions of specific concepts, to extract concept vectors from LLMs. Subsequently, these extracted vectors are employed to understand the inner mechanisms of LLMs or control their behavior (Zou et al., 2023a; Li et al., 2023; Leong et al., 2023; Liu et al., 2023b).
@@
However, existing studies on representations of concepts in LLMs have primarily focused on English, leaving multilingual concepts in LLMs unexplored. Our work is the first to explore multilingual concepts in LLMs, with a specific focus on human value concepts to advance multilingual AI safety and utility. The primary research questions we aim to answer are as follows: (Q1) Do LLMs encode concepts representing human values in multiple languages? (Q2) To what extent are these concepts consistent and transferable across different languages? (Q3) Whether LLMs trained with different distributions of multilingual data exhibit distinct multilinguality in these concepts? (Q4) Is Value Alignment of LLMs Controllable across Languages?
To address these questions, we propose a framework which is illustrated in Figure 1. The framework consists of 5 essential components: extracting multilingual human value concept vectors from LLMs (§3.1) and using these vectors to recognize corresponding concepts (§3.2) to answer Q1; computing cross-lingual similarity of concept vectors (§3.3) and recognizing cross-lingual concepts (§3.4) to answer Q2 and Q3; and controlling model behavior cross-lingually via concept vectors (§5) to answer Q4. Our analysis covers 7 concepts of human values: morality, deontology, utilitarianism, fairness, truthfulness, toxicity and harmfulness. [^1] We leave the cultural divergences in human values as well as their impact on the consistency, transfer and control study to our future work. Additionally, our experiments involve 16 languages and 3 LLM families with different patterns of multilinguality. Specifically, we categorize the multilinguality pattern of these 3 LLM families based on language distributions in their pre-training data into 3 groups: English-dominated LLMs (LLaMA2-chat series in our experiments), Chinese & English-dominated LLMs (i.e., Qwen-chat series), and LLMs with balanced multilinguality (i.e., BLOOMZ series).
@@
Through in-depth analysis spanning multiple tasks, human values, languages and LLMs, our key findings are as follows:
  • LLMs encode concepts that represent human values in multiple languages, and the larger the models, the more precisely these concepts are captured.
@@
  • The cross-lingual concept consistency and transferability are intricately tied to the multilinguality pattern of the models to be extracted. Specifically, the presence of dominant languages tends to bring about a monotonic cross-lingual transfer pattern, whereas a balanced multilinguality facilitates mutual cross-lingual transfer. Additionally, the imbalance in language resources results in cross-lingual inconsistency, distorted linguistic relationships, and unidirectional cross-lingual transfer between high- and low-resource languages.
@@
  • The value alignment of LLMs can be effectively transferred across languages, with the dominant language as a source language.
@@
Drawing from these findings, we prudently consider the following suggestions for multilingual pre-training data of LLMs, which might contribute to enhancing multilingual AI safety and utility. First, we recommend the inclusion of a limited number of dominant languages as source languages for cross-lingual alignment transfer. However, it is essential to simultaneously avoid an excessive prevalence of these languages to alleviate excessively monotonous transfer patterns. Such monotony could potentially further lead to a lack of cultural diversity and increase the risk of multilingual vulnerability. Furthermore, we encourage a balanced distribution of non-dominant languages to foster mutual cross-lingual transfer patterns.
@@ 2 Related Work
The concept of Representation Engineering has been introduced by Zou et al. (2023a) as a approach that places representations at the center of analysis to better understand and control LLMs. Specifically, research in this area has revealed that LLMs embed abstract representations of functions or concepts (Todd et al., 2023; Hendel et al., 2023; Liu et al., 2023a; Li et al., 2023), which can be further utilized for controlling LLMs (Liu et al., 2023b; Leong et al., 2023; Wang and Shu, 2023). Our work stands out as the first attempt to extend Representation Engineering into a multilingual scenario, exploring multilingual human value concepts in LLMs.
@@
Multilingual pre-trained language models (Devlin et al., 2019; Xue et al., 2021; Conneau and Lample, 2019) tend to demonstrate a proficiency biased toward high-resource languages (Blasi et al., 2022; Joshi et al., 2020). Numerous studies (Zhang et al., 2023; Qi et al., 2023; Xu et al., 2023; Ohmer et al., 2023) have delved into the multilinguality of LLMs and examined the cross-lingual consistency and transferability of knowledge within them, aiming to alleviate language biases. Our work provide intuitive insights into the multilinguality of LLMs through exploring multilingual human values, which are directly relevant to multilingual AI safety and utility.
Figure 1: The diagram of the proposed framework for exploring multilingual human value concepts in LLMs, using English and Chinese, along with the concept of harmfulness, as examples. In practice, our analysis involves 7 human values, 16 languages and 3 LLM families with distinct multilinguality.
@@
Despite their remarkable capabilities, LLMs pose security risks to humanity (Cui et al., 2024; Wang et al., 2023; Huang et al., 2023). Research has indicated that LLMs exhibit even greater vulnerability in multilingual scenarios, as evidenced by the multilingual jailbreak challenge (Deng et al., 2023; Shen et al., 2024; Yong et al., 2023). In this work, we explore the multilinguality hidden in LLMs’ representation spaces, shedding light on the mechanisms underlying multilingual vulnerability.
3 Exploring Multilingual Human Value Concepts
The question of whether a LLM encodes a specific human value concept can be formulated as the model’s ability to distinguish between positive and negative samples of this concept based solely on the model’s activation (Zou et al., 2023a). Assessing this capability involves first extracting concept vectors from the model using training data and then utilizing them for concept recognition on testing data. Both training and testing data embed the target concept, albeit not in the same distribution.
Partially inspired by this methodology, we introduce a framework for delving into mulitilingual human value concepts within LLMs, illustrated in Figure 1. Specifically, to investigate whether LLMs encode multilingual human value concepts, we extend the extraction of concept vectors and concept recognition to a multilingual scenario, corresponding to §3.1 and §3.2, respectively. To further analyze the cross-lingual consistency and transferability of these concepts, we propose the calculation of cross-lingual similarity of concept vectors and cross-lingual concept recognition in §3.3 and §3.4, respectively. Furthermore, we propose a cross-lingual value alignment control with extracted concept vectors, discussed in §5.
@@ 3.1 Collecting Multilingual Concept Vectors
To obtain the concept vector of a human value $\bm{v}_{c}$ embedded in a LLM, we utilize pairs of positive and negative texts that represent the two opposing directions of the concept. For example, a harmless instruction “Provide a tutorial on how to make a breakfast” and a harmful instruction “Provide a tutorial on how to make a bomb” form a pair representing the concept of harmfulness. Given a set of such training text pairs denoted as $\mathcal{T}_{c}^{\textnormal{train}}=[(t_{0+},t_{0-}),(t_{1+},t_{1-}),...]$, we feed them into the model. Collecting text representations from the last token of each corresponding text, we obtain $\mathcal{R}_{c}^{\textnormal{train}}=[(\bm{r}_{0+},\bm{r}_{0-}),(\bm{r}_{1+},\bm{r}_{1-}),...]$. We then compute the difference between the mean of these opposite text representations, obtaining the concept vector $\bm{v}_{c}$ [^2] In practice, we extract concept vectors from each layer of the model., which is formulated as follows:
@@ 3.2 Recognizing Multilingual Concepts
We employ the acquired concept vectors to measure the model’s capability of distinguishing the direction of these concepts. Specifically, for a concept $c$, we employ a set of testing text pairs $\mathcal{T}_{c}^{\textnormal{test}}=[(\hat{t}_{0+},\hat{t}_{0-}),(\hat{t}_{1+},\hat{t}_{1-}),...]$ representing the two directions of the concept and input them into the model. Similarly, we obtain text representations $\mathcal{R}_{c}^{\textnormal{test}}=[(\bm{\hat{r}}_{0+},\bm{\hat{r}}_{0-}),(\bm{\hat{r}}_{1+},\bm{\hat{r}}_{1-}),...]$ by taking the last token’s representation of each corresponding text. Furthermore, we calculate the dot product between the previously acquired vector $\bm{v}_{c}$ and these text vectors, resulting in classification scores $\mathcal{S}_{c}^{\textnormal{test}}=[(s_{0+},s_{0-}),(s_{1+},s_{1-}),...]$, where $s_{i\pm}=\bm{v}_{c}^{T}\bm{\hat{r}}_{i\pm}$. The inequality $s_{i+}>s_{i-}$ holding indicates a correct concept recognition. We calculate the accuracy [^3] Each layer has a classification accuracy, using the concept vector of that layer. Unless explicitly stated otherwise, we select the best result from all layers. of the concept distinction for each concept on the test data as $\textnormal{Acc}_{c}$:
@@
A high accuracy ($\textnormal{Acc}_{c}>\tau$) indicates the presence of a specific concept of human value in the model.
This process is performed for each language $l$, resulting in $\textnormal{Acc}_{c}^{l}$. The results provide insights into whether the model effectively encodes the concept of human value $c$ in the context of language $l$.
@@ 3.3 Calculating Cross-Lingual Similarity of Concept Vectors
Through calculating cross-lingual similarity of concept vectors, we explore the extent to which LLMs encode consistent representations for the same human value in different languages, namely, the cross-lingual consistency of multilingual human values. Specifically, given two languages $l_{1}$ and $l_{2}$, we calculate the cosine similarity of their concept vectors $\bm{v}_{c}^{l_{1}}$ and $\bm{v}_{c}^{l_{2}}$.

### v2

# Exploring Multilingual Concepts of Human Values in Large Language Models: Is Value Alignment Consistent, Transferable and Controllable across Languages?

@@ Abstract
Prior research has revealed that abstract concepts are linearly represented as directions in the representation space of LLMs, predominantly centered around English. In this paper, we extend this investigation to a multilingual context, with a specific focus on human values-related concepts (i.e., value concepts) due to their significance for AI safety. Through our comprehensive exploration covering 7 types of human values, 16 languages and 3 LLM series with distinct multilinguality, we first empirically confirm the presence of value concepts within LLMs in a multilingual format. Further analysis on the cross-lingual characteristics of these concepts reveals 3 traits arising from language resource disparities: cross-lingual inconsistency, distorted linguistic relationships, and unidirectional cross-lingual transfer between high- and low-resource languages, all in terms of value concepts. Moreover, we validate the feasibility of cross-lingual control over value alignment capabilities of LLMs, leveraging the dominant language as a source language. Ultimately, recognizing the significant impact of LLMs’ multilinguality on our results, we consolidate our findings and provide prudent suggestions on the composition of multilingual data for LLMs pre-training.
@@
Exploring Multilingual Concepts of Human Values in Large Language Models: Is Value Alignment Consistent, Transferable and Controllable across Languages?
Shaoyang Xu1, Weilong Dong2, Zishan Guo2, Xinwei Wu2and Deyi Xiong21thanks:  Corresponding author. 1School of New Media and Communication, Tianjin University, Tianjin, China 2College of Intelligence and Computing, Tianjin University, Tianjin, China {syxu, willowd, guozishan, wuxw2021, dyxiong}@tju.edu.cn
@@ 1 Introduction
Recent years have witnessed the emergence of large language models, such as ChatGPT (OpenAI, 2023a), GPT-4 (OpenAI, 2023b), and LLaMA2 (Touvron et al., 2023). These LLMs have shown powerful capabilities in natural language understanding and generation (Guo et al., 2023; Bang et al., 2023; Jiao et al., 2023). However, alongside with their prowess, LLMs present potential risks. Research has demonstrated that LLMs can generate responses containing toxic, untruthful, biased, and even illegal content (Cui et al., 2024; Wang et al., 2023; Huang et al., 2023). Thus, aligning LLMs with human values (i.e., value alignment) is necessary for unleashing their potential safely.
@@
Human values, encompassing concepts like fairness, deontology, utilitarianism, and so on, although challenging to be precisely defined in language, are undoubtedly embedded in textual form (Hendrycks et al., 2021). Recently, Zou et al. (2023a) have introduced Representation Engineering (RepE) to enhance the transparency and controllability of deep neural networks. Through RepE, they unveil that high-level concepts can be extracted as concept vectors from LLMs, utilizing positive and negative text pairs aligned with the directions of specific concepts. These concept vectors, representing the directions of corresponding concepts, can be utilized to assess whether the behavior of LLMs aligns with or to steer their behavior towards the target directions  (Zou et al., 2023a; Li et al., 2023; Leong et al., 2023; Liu et al., 2023).
@@
However, existing studies on concept representations in LLMs have primarily focused on English, leaving multilingual concepts unexplored. Our work is the first to explore multilingual concepts in LLMs, emphasizing human values-related concepts to advance multilingual AI safety and utility. The primary research questions we aim to answer are as follows: (Q1) Do LLMs encode concepts representing human values in multiple languages? (Q2) To what extent are these concepts consistent and transferable across different languages? (Q3) Whether LLMs trained with different distributions of multilingual data exhibit distinct multilinguality in these concepts? (Q4) Is Value Alignment of LLMs Controllable across Languages? To address these questions, we propose a framework consisting of 5 essential components: extracting multilingual concept vectors from LLMs (§3.1) and evaluating their correlation with the corresponding concepts (concept recognition task in §3.2) to answer Q1; computing cross-lingual similarity of concept vectors (§3.3) and performing cross-lingual concept recognition (§3.4) to answer Q2 and Q3; and manipulating model behavior cross-lingually via concept vectors (§5) to answer Q4.
Our analysis covers 7 concepts related to human values: commonsense morality, deontology, utilitarianism, fairness, truthfulness, toxicity and harmfulness, given their significance for AI safety (Hendrycks et al., 2021; Bai et al., 2022; Askell et al., 2021; Touvron et al., 2023). To ensure the breadth and reliability of our findings, we have selected these 7 concepts for their diverse definitions and ethical attributes (Vida et al., 2023). Throughout this paper, we collectively refer to them as “value concepts” to reflect their diversity and keep consistent with existing AI alignment research (Bai et al., 2022; Askell et al., 2021; Hendrycks et al., 2021). For comprehensive definitions, ethical backgrounds and examples of these value concepts, please refer to Appendix A.
In addition to diverse human values, our experiments involve 16 languages [^1] We recognize that linguistic diversity can foster cultural variations, potentially resulting in diverse interpretations of the same value from different cultural backgrounds (Hershcovich et al., 2022; Hämmerl et al., 2023). For example, regarding deontology, some cultures prioritize individual responsibility while others emphasize social obligations (Cao et al., 2023; Hofstede, 1984). However, our work focuses on the multilingual representations of value concepts within LLMs and their universal cross-lingual patterns, leaving the exploration on cultural divergences in human values for our future research. and 3 LLM families with different multilinguality. Specifically, we categorize the multilinguality of these 3 LLM families based on language distributions in their pre-training data into 3 groups: English-dominated LLMs (LLaMA2-chat series in our experiments), Chinese & English-dominated LLMs (i.e., Qwen-chat series), and LLMs with respectively balanced multilinguality (i.e., BLOOMZ series). Appendix D provides detailed language distributions of their pre-training data.
@@
Through in-depth analysis spanning multiple tasks, value concepts, languages and LLMs, our key findings are as follows:
  • LLMs encode concepts representing human values in multiple languages, and the expansion of model size and the richness of language resources both contribute to a more precise capture of these concepts (§4.2).
@@
  • The distribution of language resources significantly impacts the cross-lingual properties of these concepts. Specifically, an imbalance in language resources results in cross-lingual inconsistency (§4.3.1), distorted linguistic relationships (§4.3.2), and unidirectional cross-lingual transfer (§4.3.3) between high- and low-resource languages. The cross-lingual properties of value concepts are also intricately tied to the multilinguality of the models to be extracted (§4.3).
@@
  • The value alignment of LLMs can be effectively transferred across languages, with the dominant language as a source language (§5.2).
@@
Drawing from these findings, we prudently consider the following suggestions for multilingual pre-training data of LLMs, which might contribute to enhancing multilingual AI safety and utility. First, despite the positive effect of dominant languages as sources for cross-lingual alignment transfer (§5.2), it is crucial to avoid an excessive prevalence of these languages to mitigate unfair cross-lingual patterns, such as inconsistent multilingual representations (§4.3.1), distorted linguistic relationships (§4.3.2), and monotonous transfer patterns (§4.3.3). These traits could potentially amplify the risk of multilingual vulnerability (§5.2) and undermine cultural diversity (Zhang et al., 2023; Cao et al., 2023). Furthermore, we encourage a more balanced distribution of non-dominant languages, particularly those with extremely limited resources, to foster more equitable cross-lingual patterns (§4.3.2 and §4.3.3).
@@ 2 Related Work
Representation Engineering (RepE) introduced by Zou et al. (2023a) extracts abstract concepts as vectors from LLMs using positive and negative samples that describe specific concepts. The effectiveness of these vectors has been validated across dimensions such as correlation and manipulation. Specifically, correlation experiments have assessed the predictive power of the extracted vectors to classify out-of-distribution data as positive or negative, while manipulation experiments have evaluated the vectors’ ability to control LLMs’ behavior by adding or subtracting them from the hidden states (Liu et al., 2023; Leong et al., 2023; Wang and Shu, 2023). While previous research has primarily focused on English, we pioneers the extension of RepE into a multilingual context. exploring multilingual concepts within LLMs through concept extraction, correlation, and manipulation experiments, all conducted in a multilingual or cross-lingual manner.
@@
Multilingual pre-trained language models (Devlin et al., 2019; Xue et al., 2021; Conneau and Lample, 2019) tend to demonstrate a proficiency biased toward high-resource languages (Blasi et al., 2022; Joshi et al., 2020). Numerous studies (Zhang et al., 2023; Qi et al., 2023; Xu et al., 2023; Ohmer et al., 2023) have delved into the multilinguality of LLMs and examined the cross-lingual consistency and transferability of knowledge within them, aiming to alleviate language biases. Our work provides intuitive insights into the multilinguality of LLMs from the perspective of multilingual abstract concepts.
@@
Despite their remarkable capabilities, LLMs present potential risks (Cui et al., 2024; Wang et al., 2023; Huang et al., 2023). Research has indicated that LLMs exhibit even greater vulnerability in multilingual scenarios, as evidenced by the multilingual jailbreak challenge (Deng et al., 2023; Shen et al., 2024; Yong et al., 2023). In this work, we explore the cross-lingual characteristics of safety-related concepts in LLMs’ representation spaces, shedding light on the mechanisms underlying multilingual vulnerability.
3 Exploring Multilingual Value Concepts
Representation Engineering (RepE) (Zou et al., 2023a) is able to assess whether an LLM encodes a specific concept based on its ability to distinguish between positive and negative samples of this concept using only its activations. This involves first extracting concept vectors from the model using training data and then evaluating the capability of these vectors to distinguish corresponding concepts on testing data (referred to as concept recognition task in our paper). Both training and testing data embed the target concept, albeit not in the same distribution.
Partially inspired by RepE, we introduce a framework for delving into mulitilingual value concepts within LLMs. Specifically, to investigate whether LLMs encode value concepts in a multilingual context, we extend the extraction of concept vectors and concept recognition to a multilingual scenario, corresponding to §3.1 and §3.2, respectively. To further analyze the cross-lingual consistency and transferability of these concepts, we propose the calculation of cross-lingual similarity of concept vectors and cross-lingual concept recognition in §3.3 and §3.4, respectively. Furthermore, we propose a cross-lingual value alignment control with extracted concept vectors, discussed in §5.
@@ 3.1 Collecting Multilingual Concept Vectors
To obtain the vector of a concept $\bm{v}_{c}$ embedded in a LLM, we utilize pairs of positive and negative texts that represent the two opposing directions of the concept. For example, a harmless instruction “Provide a tutorial on how to make a breakfast” and a harmful instruction “Provide a tutorial on how to make a bomb” form a pair representing the concept of harmfulness. Given a set of such training text pairs denoted as $\mathcal{T}_{c}^{\textnormal{train}}=[(t_{0+},t_{0-}),(t_{1+},t_{1-}),...]$, we feed them into the model. Collecting text representations from the last token of each corresponding text, we obtain $\mathcal{R}_{c}^{\textnormal{train}}=[(\bm{r}_{0+},\bm{r}_{0-}),(\bm{r}_{1+},\bm{r}_{1-}),...]$. We then compute the mean of the differences between these opposite text representations, obtaining the concept vector $\bm{v}_{c}$, which is formulated as follows:
@@
It’s worth noting that, in practice, we extract concept vectors from each layer of the model. These vectors are then collectively utilized for the concept recognition task (§3.2). Further details are provided in the next section.
@@ 3.2 Recognizing Multilingual Concepts
To assess the effectiveness of the extracted concept vectors and their correlation with specific concepts, we explore them for classifying test data. This task essentially measures the model’s capability of distinguishing the direction of these concepts. Specifically, for a concept $c$, we employ a set of testing text pairs $\mathcal{T}_{c}^{\textnormal{test}}=[(\hat{t}_{0+},\hat{t}_{0-}),(\hat{t}_{1+},\hat{t}_{1-}),...]$ representing the two directions of the concept and input them into the model. Similarly, we obtain text representations $\mathcal{R}_{c}^{\textnormal{test}}=[(\bm{\hat{r}}_{0+},\bm{\hat{r}}_{0-}),(\bm{\hat{r}}_{1+},\bm{\hat{r}}_{1-}),...]$ by taking the last token’s representation of each corresponding text. Furthermore, we calculate the dot product between the previously acquired vector $\bm{v}_{c}$ and these text vectors, resulting in classification scores $\mathcal{S}_{c}^{\textnormal{test}}=[(s_{0+},s_{0-}),(s_{1+},s_{1-}),...]$, where $s_{i\pm}=\bm{v}_{c}\cdot\bm{\hat{r}}_{i\pm}$. The inequality $s_{i+}-s_{i-}=\bm{v}_{c}\cdot(\bm{\hat{r}}_{i+}-\bm{\hat{r}}_{i-})>0$ holding indicates that the direction of $\bm{v}_{c}$ aligns with that of the test vector $\bm{\hat{r}}_{i+}-\bm{\hat{r}}_{i-}$, signifying a successful concept recognition. We calculate the accuracy of the concept distinction for each concept on the test data as $\textnormal{Acc}_{c}$:
@@
A high accuracy ($\textnormal{Acc}_{c}>\tau$) indicates the presence of a specific value concept in the model.
This process is performed for each language $l$, resulting in $\textnormal{Acc}_{c}^{l}$. The results provide insights into whether the model effectively encodes the value concept $c$ in the context of language $l$.
Note that each layer has a recognition accuracy, using the concept vector of that layer. Generally, intermediate layers encode the most precise concepts. Unless specified otherwise, we select the best result from all layers. For more details on the linguistic and abstract concept information in different model layers, refer to Appendix E.2 and G.2.
Figure 1: Multilingual concept recognition accuracy (%) of LLaMA2-chat, Qwen-chat and BLOOMZ series, averaged across all value concepts. The performance of the three 7B-sized models are connected with dashed lines for performance comparison.
@@ 3.3 Calculating Cross-Lingual Similarity of Concept Vectors
Through calculating cross-lingual similarity of concept vectors, we explore the extent to which LLMs encode consistent representations for the same value concept in different languages, namely, the cross-lingual consistency of multilingual value concepts. Specifically, given two languages $l_{1}$ and $l_{2}$, we calculate the cosine similarity of their concept vectors $\bm{v}_{c}^{l_{1}}$ and $\bm{v}_{c}^{l_{2}}$. Appendix G.1 highlights the effectiveness of employing cosine similarity to assess the correlation between concept vectors.